OpenStudy (help_people):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n? n ≥ - the quantity 4 times c minus 3 all over p n ≥ - the quantity 4 times c minus 13 all over p n ≤ - the quantity 4 times c minus 3 all over p n ≤ - the quantity 4 times c minus 13 all over p

OpenStudy (help_people):

@welshfella

Nnesha (nnesha):

first of all distribute parentheses by outside term

OpenStudy (help_people):

so -np-5<=4x-8

OpenStudy (amoodarya):

-np - 5 ≤ 4(c - 2) np+5

Nnesha (nnesha):

alright yes that's right you have to isolate n so move -5 to the right side and remember when you MULTIPLY or DIVIDE by negative number then you should flip the inequality sign

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @help_people so -np-5<=4x-8 \(\color{blue}{\text{End of Quote}}\) 4c*

OpenStudy (amoodarya):

np +5 >=- 4(c - 2) np >= -4c +8-5 np> = -4c+3

OpenStudy (help_people):

ok

OpenStudy (help_people):

why do i do after?

OpenStudy (amoodarya):

divide by p

OpenStudy (help_people):

np/p=-4c+3/p n>=4c-3/p

OpenStudy (anonymous):

-np - 5 ≤ 4(c - 2) -np - 5 ≤ 4c - 8 adding +5 for both sides -np=<4c-8+5 -np=< 4c-3 multiply by (-1) np >=-4c+3

OpenStudy (help_people):

based on that a is the answer?

Nnesha (nnesha):

based on what ? \[\rm n \ge \frac{ -4c \color{reD}{+}3 }{ p }\] n ≥ - the quantity 4 times c **minus** 3 all over p

Nnesha (nnesha):

or you can write it as \[\rm n \ge \frac{ 3 -4c }{ p }\] but still 3 is positive take a screenshot

OpenStudy (help_people):

well none of this options have that so a right?

Nnesha (nnesha):

a is incorrect but bec none of options are right so yeah a.......:/

Nnesha (nnesha):

n ≥ - the quantity 4 times c minus 3 all over p \[\rm n \ge \frac{ -4c \color{reD}{-}3 }{ p }\] \[\rm n \ge \frac{ \color{reD}{-} (4c \color{reD}{+}3) }{ p }\] negative sign is common so take it out mhm

Nnesha (nnesha):

if we do backwards then that^^ makes sense

OpenStudy (help_people):

ok just to claify because i do not want to get it wrong a is correct

OpenStudy (help_people):

@nn

Nnesha (nnesha):

i guess YES

OpenStudy (help_people):

ok may you help w. another ?

Nnesha (nnesha):

i'll try

OpenStudy (help_people):

ok

OpenStudy (help_people):

Which equation is not equivalent to the formula e = mc? m equals e over c c equals e over m e = cm m equals c over e

OpenStudy (help_people):

i think its d @Nnesha

Nnesha (nnesha):

how ? can you explain a littlebit ?

Nnesha (nnesha):

:-)

OpenStudy (help_people):

well sure i solve for m and c idk about the e=cm but that leaves us with m=c/e

Nnesha (nnesha):

solve for e m =c/e

Nnesha (nnesha):

yes that's right m = c/e equal to em =c which is not equivalent to e =mc

OpenStudy (help_people):

wait so d is right??

OpenStudy (help_people):

it is not equally to e=mc because what u said before confused me :D

Nnesha (nnesha):

ye d is right

Nnesha (nnesha):

equation is e = mc m TIMES c and D is m = c/e which is em = c solve for e you will get e = m/c m OVER c